OpenStudy (anonymous):

Quick Question: What are the steps in solving for Y for this equation: X=720+140y/y ? thanks so much!

OpenStudy (anonymous):

140y/y = 140 ?

OpenStudy (anonymous):

The answer is Y = 720/x-140 , but I don't know how to get to that answer.

OpenStudy (campbell_st):

multiply both sides by y... to remove the denominator in the fraction \[Xy = 720 + 140y\] next subtract 140y from both sides \[Xy - 140y = 720\] take the common factor out on the left hand side \[y(X - 140) = 720\] hope this helps..

OpenStudy (anonymous):

Thank you so much!! Been racking my brain for the last half hour!